HUBBLE SPACE TELESCOPE DAILY REPORT #3034 PERIOD COVERED: 0000Z (UTC) 01/14/02 - 0000Z (UTC) 01/15/02 Daily Status Report as of 015/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.2 Completed Two Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.3 Completed WF/PC-2 9060 (Photometry of a Statistically Significant Sample of Kuiper Belt Objects) The WF/PC-2 was used to propel the physical study of KBOs forward by performing accurate photometry at V, R, and I on a sample of up to 150 KBOs. The sample is made up of objects that will be observed at thermal infrared wavelengths by SIRTF and will be used with those data to derive the first accurate diameters and albedos for a large sample of KBOs. The observations completed nominally. 1.4 Completed Eight Sets of WF/PC-2 8942 (Cycle 10 Intflat Sweeps and Linearity Test) The WF/PC-2 was used to monitor the pixel-to-pixel flatfield response and provide a linearity check. These intflat sequences are done once during the year and the images will provide a backup database in the event of complete failure of the visflat lamp as well as allow monitoring of the gain ratios. There were no reported problems. 1.5 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.6 Completed Six Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.7 Completed STIS/MA2 8573 (Newborn Planets and Brown Dwarf Companions in IC 348) The Space Telescope Imaging Spectrograph (MA2) was used to search for young giant planets and brown dwarfs around ~100 low-mass stars and brown dwarfs in the nearby cluster IC 348. The observations completed with no reported problems. 1.8 Completed STIS/CCD/MA1 9067 (UV Detectability of Bright Quasars in the Sloan Fields) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to take MAMA spectra of approximately 30 new, bright, high-redshift quasars in each of the next three cycles. The observations completed with no reported problems 1.9 Completed Four Sets of WF/PC-2 8935 (Cycle 10 Standard Darks) The WF/PC-2 was used to obtain dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs. The proposal completed with no reported problems. 1.10 Completed Three Sets of STIS/CCD/MA2 8471 (D/H In Quasar Absorption Line Systems) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to record the high Lyman series lines of two damped Lyman-alpha systems in the coverage of the near-UV band, where the interference from randomly situated Lyman-alpha forest lines is much lower than at visible wavelengths. A secondary objective of this program is to observe various atomic lines shortward of Lyman-alpha in these systems {such as absorption from the O VI doublet arising from very hot gas}, and to register absorption features from the Lyman and Werner bands of molecular hydrogen if the molecular abundances are high enough to make them detectable. The proposal completed with no reported problems. 1.11 Completed Two Sets of STIS/MA1/MA2 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.12 Completed WF/PC/STIS/CCD-2 9125 (Are there Large Dust Grains in Orion's Proto-Planetary Disks?) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD)were used to acquire dithered PC images of six of the largest circumstellar disks in the Orion nebula to study the unique properties of dust within these systems. No anomalous activity occurred 1.13 Completed WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. No problems were encountered. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 8 Successful: 8 2.2 FHST Updates: Scheduled: 13 Successful: 13 2.3 Operations Notes: Using ROP SR-1A, the SSR-1 EDAC error counter was cleared twice. Operations personnel are supporting Joint Integrated Simulation #4 (EVA-3/Planning/EVA-4). The simulation began at 9:00 a.m. yesterday and will continue until approximately 8:00 p.m. this evening. A TTR was written for a required re-transmit at 014/130103Z during a NSSC-1 load. Accordingly, ROP NS-5 was used to reset the SI C&DH error. The ephemeris table was uplinked at 015/0112Z per ROP DF-07A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW